Sheridan Corporation issued 368 shares of $10 par value common stock and 123 shares of $50 par value preferred stock for a lump sum of $16,587. The common stock has a market price of $20 per share, and the preferred stock has a market price of $90 per share.

Prepare the journal entry to record the issuance. (Round intermediate calculations to 6 decimal places, e.g. 0.546872 and final answers to 0 decimal places, e.g., 1,520. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit
Answer #1

Ans:

Account Title & Explanation

Debit

Credit

Cash

$16,587

Common stock (368*10)

$3680

Preferred stock ( 123*50)

$6150

Additional Paid in Capital - Common Stock (6624-3680)

$2944

Additional Paid in Capital - Preferred Stock (9963-6150)

$3813

Working

Total market value of 319 share of common stock = 368*20 = 7360

Total market value of 119 share of Preferred stock = 123*90 = 11070

Total Market Value = $ 18430

lump sum of $16,587 is distributed to common stock & preferred stock in ratio of market value

Proportionate Value of Common Stock = 16587*7360/18430 = 6624

Proportionate Value of Preferred Stock = 16587*11070/18430= 9963